# Hao Liu1    Yu Hu1    Rakiba Rayhana1    Ling Bai1&Zheng Liu1

### v1

@@ Abstract
Bed-related falls remain a leading source of injury in hospitals and long-term-care facilities, yet many commercial alarms trigger only after a patient has already left the bed. We show that early bed-exit intent can be predicted using only four low-cost load cells mounted under the bed legs. The resulting load signals are first converted into a compact set of complementary images: an RGB line plot that preserves raw waveforms and three texture maps—recurrence plot, Markov transition field, and Gramian angular field—that expose higher-order dynamics. We introduce ViFusionTST, a dual-stream Swin Transformer that processes the line plot and texture maps in parallel and fuses them through cross-attention to learn data-driven modality weights.
To provide a realistic benchmark, we collected six months of continuous data from 95 beds in a long-term-care facility. On this real-world dataset ViFusionTST reaches an accuracy of 0.885 and an F1 score of 0.794, surpassing recent 1D and 2D time-series baselines across F1, recall, accuracy, and AUPRC. The results demonstrate that image-based fusion of load-sensor signals for time series classification is a practical and effective solution for real-time, privacy-preserving fall prevention.
@@ 1 Introduction
Traditionally, 1D sequence models—ranging from multi-layer perceptrons (MLPs) to deep convolutional neural networks (CNNs)/recurrent neural networks (RNNs)—were developed expressly for time-series signals such as bed loads Zeng et al. (2023); Bai et al. (2018). In recent years, the field has progressed to Time Series Transformers Zerveas et al. (2021) and even large language models (LLMs) Jin and others (2024). Meanwhile, Vision Transformers Dosovitskiy and others (2021); Liu et al. (2021) have achieved remarkable success in computer vision, motivating researchers to explore whether the same architectures can model temporal data Chen et al. (2024). This exploration is enabled by a rich set of techniques that turn 1D time series into 2D images: line plots for multivariate signals and, for univariate signals, recurrence plots (RP), Markov transition fields (MTF), and Gramian angular fields (GAF) Li et al. (2024); Costa et al. (2024). Compared with raw numeric sequences, these images give both humans and neural networks a more intuitive view of temporal patterns. Vision methods for time series bring two practical benefits: 1) by pretraining on a large number of images, vision models have already learned common patterns such as trends, cycles, and peaks; 2) representing a long sequence as a compact image reduces input length, easing the burden of modeling long-range dependencies Ni et al. (2025).
@@ 2.1 Bed Exit Prediction
Predicting a bed exit before it happens remains difficult in healthcare practice. Most existing systems only raise an alarm after the patient leaves the bed Lin et al. (2022); Chang et al. (2021); Inoue and Taguchi (2020), giving caregivers little time to act.
Vision-based research has taken the first steps toward prediction. Chen et al. Chen et al. (2018) used a depth camera and a 3D CNN to detect two pre-exit motions: torso rotation while sitting and rolling out of bed. However, their study provided only qualitative examples of predictive capability without a quantitative analysis of its accuracy. BED-net Bu et al. (2021) extended this idea with an R(2+1)D backbone and reached 87.7% accuracy detecting transitional movements, yet performance dropped when the camera angle changed or blankets covered the subject.
@@
All existing methods either compromise patient privacy Ansari et al. (2022); Wang et al. (2022), require expensive hardware, or cover only a limited set of bed-exit behaviors. We present the first bed-exit prediction system that works with a single low-cost load cell mounted under one bed leg. The sensor fits any bed frame, adds minimal hardware cost, and scales easily. Moreover, our model learns the common cues across diverse bed-exit motions directly from data, without hard-coding specific transitions, making it flexible for real-world use.
@@ 3.1 Bed Monitoring Platform Overview
In this study, we developed a bed monitoring platform that enables unobtrusive, non‐contact monitoring of hospital beds, as illustrated in Fig. 1. The bed sensor employs a strain gauge–based load cell housed in an enclosure with a built‐in ramp and platform, allowing a bed wheel to roll up and remain securely on top. The sensor is powered by a standard power cord and transmits data wirelessly to our AI platform through an IoT (Internet of Things) network.
@@ 3.2.4 Gramian Angular Field (GAF)
GAF Wang et al. (2015); Wang and Oates (2015) represents time-series values in polar coordinates and captures pairwise angular relationships. First, we rescale the series $x_{t}$ to $[-1,1]$, obtaining $\tilde{x}_{t}$. Then encode each $\tilde{x}_{t}$ as an angle
@@ 4.3.1 Effect of Fusion Method
Feature Concatenation 0.658 0.582 0.756 0.838 0.761
Cross-Attention Fusion (Ours) 0.794 0.827 0.763 0.885 0.868
% Improvement +20.7% +42.1% +0.9% +5.6% +14.1%
Table 2: Evaluation of fusion methods.

### v4

@@ Abstract
Bed-related falls remain a major source of injury in hospitals and long-term care facilities, yet many commercial alarms trigger only after a patient has already left the bed. We show that early bed-exit intent can be predicted using only one low-cost load cell mounted under a bed leg. The resulting load signals are first converted into a compact set of complementary images: an RGB line plot that preserves raw waveforms and three texture maps—recurrence plot, Markov transition field, and Gramian angular field—that expose higher-order dynamics. We introduce ViFusionTST, a dual-stream Swin Transformer that processes the line plot and texture maps in parallel and fuses them through cross-attention to learn data-driven modality weights. To provide a realistic benchmark, we collected six months of continuous data from 95 beds in a long-term-care facility. On this real-world dataset ViFusionTST reaches an accuracy of 0.885 and an F1 score of 0.794, surpassing recent 1D and 2D time-series baselines across F1, recall, accuracy, and AUPRC. The results demonstrate that image-based fusion of load-sensor signals for time series classification is a practical and effective solution for real-time, privacy-preserving fall prevention.
@@ 1 Introduction
Traditionally, 1D sequence models, ranging from multi-layer perceptrons (MLPs) to deep convolutional neural networks (CNNs)/recurrent neural networks (RNNs), were developed expressly for time-series signals such as bed loads Zeng et al. (2023); Bai et al. (2018). In recent years, the field has progressed to Time Series Transformers Zerveas et al. (2021) and even large language models (LLMs) Jin and others (2024). Meanwhile, Vision Transformers Dosovitskiy and others (2021); Liu et al. (2021) have achieved remarkable success in computer vision, motivating researchers to explore whether the same architectures can model temporal data Chen et al. (2024). This exploration is enabled by a rich set of techniques that turn 1D time series into 2D images: line plots for multivariate signals and, for univariate signals, recurrence plots (RP), Markov transition fields (MTF), and Gramian angular fields (GAF) Li et al. (2024); Costa et al. (2024). Compared with raw numeric sequences, these images give both humans and neural networks a more intuitive view of temporal patterns. Vision methods for time series bring two practical benefits: 1) by pretraining on a large number of images, vision models have already learned common patterns such as trends, cycles, and peaks; 2) representing a long sequence as a compact image reduces input length, easing the burden of modeling long-range dependencies Ni et al. (2025).
@@ 2.1 Bed Exit Prediction
Predicting a bed exit before it happens remains difficult in healthcare practice. Most existing systems only raise an alarm after the patient leaves the bed Lin et al. (2022); Chang et al. (2021); Inoue and Taguchi (2020), leaving caregivers little time to act.
Early studies have sought to bridge this gap by employing camera-based human-action recognition methods. Chen et al. Chen et al. (2018) used a depth camera and a 3D CNN to detect two pre-exit motions: torso rotation while sitting and rolling out of bed. However, their study provided only qualitative examples of predictive capability without a quantitative analysis of its accuracy. BED-net Bu et al. (2021) extended this idea with an R(2+1)D backbone and reached 87.7% accuracy detecting transitional movements, yet performance dropped when the camera angle changed or blankets covered the subject.
@@
Across different modalities, existing methods either expose sensitive imagery of patients, require expensive hardware, or cover only a limited set of bed-exit behaviors. We present the first bed-exit prediction system that works with a single low-cost load cell mounted under one bed leg. The sensor fits any bed frame, adds minimal hardware cost, and scales easily. Moreover, our model learns the common cues across diverse bed-exit motions directly from data, without hard-coding specific transitions, making it flexible for real-world use.
@@ 3.1 Bed Monitoring Platform Overview
In this study, we develop a bed monitoring platform that enables unobtrusive, non‐contact monitoring of hospital beds, as illustrated in Fig. 1. The bed sensor employs a strain gauge–based load cell housed in an enclosure with a built‐in ramp and platform, allowing a bed wheel to roll up and remain securely on top. The sensor is powered by a standard power cord and transmits data wirelessly to our AI platform through an IoT (Internet of Things) network.
@@ 3.2.4 Gramian Angular Field (GAF)
GAF Wang et al. (2015); Wang and Oates (2015) represents time-series values in polar coordinates and captures pairwise angular relationships. First, we rescale the series $x_{t}$ to [1,1][-1,1], obtaining $\tilde{x}_{t}$. Then encode each $\tilde{x}_{t}$ as an angle
@@ 4.3.1 Effect of Fusion Method
Value Diff Value Diff Value Diff Value Diff Value Diff
Early Concatenation 0.658 -17.1% 0.582 -29.6% 0.756 -0.9% 0.838 -5.3% 0.761 -12.3%
Gated Fusion 0.722 -9.1% 0.672 -18.7% 0.780 +2.2% 0.861 -2.7% 0.814 -6.2%
Mid Concatenation 0.764 -3.8% 0.720 -12.9% 0.814 +6.7% 0.881 -0.5% 0.847 -2.4%
Cross Fusion (Ours) 0.794 ref. 0.827 ref. 0.763 ref. 0.885 ref. 0.868 ref.
Table 3: Evaluation of fusion methods. “Diff” columns indicate relative difference compared to Cross Fusion (Ours).
Vision Encoder F1 Score $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ Accuracy $\uparrow$ AUPRC $\uparrow$
Resnet-50 He et al. (2016) 0.727 0.663 0.804 0.866 0.817
MambaVision-Base Hatamizadeh and Kautz (2025) 0.731 0.665 0.812 0.869 0.813
MambaVision-Small Hatamizadeh and Kautz (2025) 0.729 0.655 0.824 0.870 0.821
MambaVision-Tiny Hatamizadeh and Kautz (2025) 0.759 0.702 0.825 0.880 0.830
SwinV2-Base Liu et al. (2022) 0.701 0.656 0.753 0.850 0.785
SwinV2-Small Liu et al. (2022) 0.734 0.815 0.668 0.842 0.782
ViT-Large Dosovitskiy and others (2021) 0.721 0.649 0.811 0.865 0.783
ViT-Base Dosovitskiy and others (2021) 0.731 0.690 0.778 0.864 0.822
Swin-Base Liu et al. (2021) 0.766 0.797 0.738 0.870 0.817
Swin-Small Liu et al. (2021) 0.775 0.783 0.767 0.878 0.822
Swin-Tiny Liu et al. (2021) 0.794 0.827 0.763 0.885 0.868
Table 4: Performance Comparison of Vision Encoders
@@ 4.3.2 Effect of Input Modalities
4.3.3 Effect of Vision Encoders
To quantify the impact of different pretrained vision backbones on bed-exit prediction, we replace our Swin-Tiny encoders with a variety of alternatives, while keeping all other components fixed. Table 4 reports the performance for each encoder. The Swin-Tiny encoder achieves the best overall performance, surpassing all other encoders, including its larger variants. Among non-Swin architectures, the MambaVision delivers the next best results, with MambaVision-Tiny achieving F1 = 0.759 and AUPRC = 0.830. This suggests that replacing self-attention layers with Mamba blocks can also effectively capture temporal image features while reducing computational cost.
Notably, within each model family, the smaller variant consistently outperforms its larger counterpart (e.g., ViT-Base vs. ViT-Large). We attribute this to the relatively low complexity of temporal image representations compared to natural images. Larger models tend to overfit on these simpler patterns, leading to degraded generalization. Overall, these findings emphasize the importance of selecting appropriately sized encoders for time-series image inputs, since larger models are not necessarily better.
4.3.4 Prediction Visualization
Figure 4 presents a representative bed-exit event, where the load signal (blue) and the predicted bed-exit probability (green) are plotted over time. During the initial stable-in-bed phase, the patient remains motionless. Consequently, the load is flat and the model outputs a probability of 0, suppressing unnecessary alarms. As the patient transitions toward bed exit, subtle posture shifts introduce high-frequency load oscillations. These precursor movements cause the probability to rise steadily. The ward’s nurse call system employs a fixed alarm threshold of 0.5 (dashed line). The model first exceeds this threshold six minutes before the true exit time. This advance notice is clinically valuable: caring staff receive a single precise alert rather than a series of premature or fluctuating warnings. These observations demonstrate that the ViFusionTST model can identify preparatory movements that simple load-drop heuristics cannot detect, while providing interpretable and actionable predictive results suitable for deployment in the real world.
Figure 4: ViFusionTST prediction probability over time for a bed-exit event.
@@ 5 Conclusion
Acknowledgments
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant ALLRP 571106-21, and conducted in collaboration with Tochtech Technologies Ltd., who provided the bed load sensor modules and online data platform for data collection.